OpenStudy (anonymous):

Reaction is CO2(g) + C(s) <=> 2CO(g) Kp = 5.7, 1200K Calcualte the total pressure at equilibrium when 4.45g CO2 is introduced into a 10.0L container and heated to 1200K in the presence of 2.00g of graphite

OpenStudy (anonymous):

What effect does the mass of graphite have on this?

OpenStudy (anonymous):

first step find the pressure exerted pv=nRT use this relation and find the pressure at initial state

OpenStudy (anonymous):

so u get the patial pressure of CO2 and graphite so this is the pressure at initial state at equilibrium, co2+ c------2co pco2-k pc-k 2k Kc=2k^2/pco2^2 *pc^2

OpenStudy (anonymous):

How can a solid have partial pressure?

OpenStudy (anonymous):

very good question!................is the graphite in pure form?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

yes it should be then concentration=moles/volume=density/molar mass density and molar mass being constant graphite has concentration=1 therefore Pc=1

OpenStudy (anonymous):

I do not understand what effect does the mass of graphite have on this question. The question after this asks me to do this entire process again but for mass of graphite = 0.50g...

OpenStudy (anonymous):

Would not the graphite burn as well?

OpenStudy (anonymous):

yes but not contribute to kc which we are bothered abt here and to equilibrium concentration

OpenStudy (anonymous):

So you are saying that in both cases, the answer is the same? (ie. when it is 0.50g graphite, or the mass given in this question)...

OpenStudy (anonymous):

yeah it souind vague but yeah thats the thing i am sure about it it is only when u have 2gms of c that the raection proceeds if it were less then half the reaction will take place wat i mewan by this is for 1 mol of co2 1 mol of c is to be present to give 2 mol of CO so thats basically the idea

OpenStudy (mos1635):

n1 =moles of CO2 (find them) n2 = moles of C (find them) CO2(g) + C(s) <=> 2CO(g) n1 n2 0 -x -x +2x n1-x n2-x 2x Kp=P^2 (CO)/Pco2 ..........use PV=nRT Kp=(2x)^2/n1-x *RT/V solve to x then P(total)=n(total)*R*T/V......Ptot=....

OpenStudy (anonymous):

Salini, you might want to review. Both answers being the same is incorrect.

OpenStudy (anonymous):

Mos, the Carbon solid is a constant. So it does not take place in an equilibrium ICE table.

OpenStudy (mos1635):

yes you are right i should put n(total) as n (gas total)=n(co2)+n(co)